OpenStudy (anonymous):

The area of a rectangle is 44 yd2 , and the length of the rectangle is 3 yd less than twice the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

let the width = w then the length is 2w - 3 so then you can find the area \[A = w(2w - 3) ~~~or~~~A = 2w^2 - 3w\] you know the area so \[44 = 2w^2 - 3w \] rewriting the equation \[2w^2 - 3w - 44 = 0\] use the general quadratic formula to solve for w, only consider the positive value for w. hope it helps... p.s. the with is a mixed number....

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

so the width is 11/2 ? @campbell_st

OpenStudy (anonymous):

and the length is 8?

OpenStudy (kendricklamar2014):

Here is how I would do it: A = L*W L = 2W - 3 A = (2W-3)*W 44 = 2W^2 - 3W 2W^2 - 3W - 44 = 0 ((3 +- sqrt((-3)^2 - 4*2*-44))/4 (3 +- sqrt( 9 + 352))/4 (3 +- 19)/4 W= -16/4 and 22/4 However, W must be positive as we are talking about a measurement. So W = 22/4 = 5.5 Plug that back into our original 44 = L * 5.5 44/5.5 = 8 So our dimensions are 8 x 5.5.

OpenStudy (anonymous):

oh ok got it! thank you!!! =)

OpenStudy (kendricklamar2014):

You're Welcome :)